# Q-Probe: A Lightweight Approach to Reward Maximization for Language Models

### v1

@@ Abstract
We present an approach called Q-probing to adapt a pre-trained language model to maximize a task-specific reward function. At a high level, Q-probing sits between heavier approaches such as finetuning and lighter approaches such as few shot prompting, but can also be combined with either. The idea is to learn a simple linear function on a model’s embedding space that can be used to reweight candidate completions. We theoretically show that this sampling procedure is equivalent to a KL-constrained maximization of the Q-probe as the number of samples increases. To train the Q-probes we consider either reward modeling or a class of novel direct policy learning objectives based on importance weighted policy gradients. With this technique, we see gains in domains with ground-truth rewards (code generation) as well as implicit rewards defined by preference data, even outperforming finetuning in data-limited regimes. Moreover, a Q-probe can be trained on top of an API since it only assumes access to sampling and embeddings. Code: https://github.com/likenneth/q_probe.
@@ 1 Introduction
On the other hand, there is evidence that the capabilities required for these downstream tasks have already been learned during pre-training, and the task of adaptation is merely extract them from the wide spectrum of pre-trained capabilities. For example,  Zaken et al. [2021] propose that extremely parameter-efficient finetuning is evidence that the finetuning process is mostly about “exposing knowledge induced by language-modeling training”, while Saunders et al. [2022] find that pre-trained language models are usually better at discriminating than generating answers.
Motivated by this line of thought, we present a lightweight approach reward maximization. For each downstream task, we keep the whole pre-trained model frozen and only train a small probe which is the same dimension as the residual stream [Alain and Bengio, 2016]. We call our method Q-probe as it “probes” the expected utility of a completion (action) given a certain prompt (state).
To leverage the Q-probe at inference to generate samples we perform a sort of rejection sampling. Specifically, we first draw $k$ sampled completions from the LLM given the input prompt and also store the embedding of each prompt-completion pair. The Q-probe then predicts a value for each embedding which determine the logits for $k$-way softmax distribution that we use to sample the chosen completion. In theory, we show that this procedure maximizes the KL-constrained value of the probe as $k$ tends to infinity.
Figure 1: An illustration of the Q-probe inference procedure. Given a prompt, we use the language model to generate $k=3$ completions (in this case programs) and the respective embeddings of the $k$ prompt-completion pairs. Then the linear Q-probe maps the features into the logits of a softmax distribution. We obtain our final sample from the Q-probe by sampling from this distribution.
First, we evaluate Q-probes with access to ground truth rewards on coding benchmarks our best Q-probe achieves $17\%$ higher accuracy on MBPP [Austin et al., 2021] compared to the base Code-LLaMA-7B [Roziere et al., 2023] and outperforms finetuning on successes with LORA [Hu et al., 2021] and few shot prompting. One key component of the results is a novel objective for training the Q-probes via direct policy learning. We find that rather than training the Q-probe to model the rewards, it is more effective to use an importance weighted policy gradient objective. Since we only need access to samples and embeddings we can train Q-probes on API-based models where gains are more modest ($3\%$ improvement over base model) due to a stronger base model and lack of access to internal model embeddings.
@@
Finally, in terms of computational cost, we should note that using a Q-probe requires substantially less training compute, but more inference-time compute when compared to finetuning. In our experiments, we can train a Q-probe in a few seconds (since it is just a 4096-dimensional linear model) whereas even parameter efficient finetuning [Hu et al., 2021] takes several hours. But, at inference, we draw $k$ samples from the base model rather than 1 from a finetuned model, although improvements in parallel and speculative decoding are making batched decoding easier [Fang et al., 2021, Yu et al., 2022]. That said, it is also worth noting again that finetuning and Q-probing are not mutually exclusive. Indeed in our preference experiments, we find that the combination yields better performance than either method alone.
@@ 2 Related work
Q-probes leverage the idea of probing to solve reward maximization problems. This idea builds on prior work that uses probe for understanding the internals of neural networks [Alain and Bengio, 2016, Belinkov, 2016, Li et al., 2022]. A probe is a classifier or regressor which takes internal activations of a network as its input and is trained to predict a feature of interest, e.g., part of speech, parse tree depth, or the expected reward in our case.
@@
Rejection sampling for reward maximization is not a new idea. In fact, Ganguli et al. [2022], Rafailov et al. [2023] also evaluate rejection sampling as one of their baselines. However, their selector model is instantiated by the preference language model trained in a similar way to the first stage of RLHF by Ouyang et al. [2022]. This version of rejection sampling not only involves higher training cost but is also double the inference cost to run the reward model while evaluating the Q-probe is essentially free in comparison to the base model.
@@
An important line of training-free adaptation methods centers around prompting [Salewski et al., 2023] which includes in-context learning (ICL,  Min et al., 2022) and Chain-of-thoughts (CoT,  Wei et al., 2022). Though it enjoys great flexibility, Mosbach et al. [2023] reveal by a closer examination that finetuning still outperform prompting methods. Prompting could also be sensitive to prompt engineering [Lu et al., 2021] and takes up valuable context window, limiting the amount of data we can feed into it for a fair comparison with Q-probe and finetuning.
@@
There are also a host of other inference-time techniques designed for coding and reasoning settings [Zhou et al., 2023a, Shinn et al., 2023, Yao et al., 2023]. However, they require access to the feedback from environment at test time which is different from the one-pass setting considered by us.
@@ 3.1 Feedback: oracle rewards and preferences
Note, there is a large literature of prior work on using reinforcement learning directly to finetune language models when given access to oracle reward functions, e.g. for single turn language tasks [Schulman et al., 2017, Snell et al., 2022, Ramamurthy et al., 2022, Chang et al., 2023] or in multiturn settings [Zhou et al., 2023b, Abdulhai et al., 2023]. In contrast, we focus on a lighter weight approach that only requires training probes, but shows how probe training can approximate traditional RL objectives.
@@
    Online. With online access we can query the reward or preference of any action $a$ or action pair $(a_{0},a_{1})$ from any context $x_{i}$ in the training set to get $r(x_{i},a)$. This setting is reasonable if we have unit tests for programming or a human in the loop for preference learning.
@@ 3.3 Access to the LLM
    Sampling from the LM distribution $p_{0}$. Given a context $x$ we can sample a completon $a$ from $p_{0}(\cdot|x)$.
@@
We do not in general assume access to the underlying model to allow for finetuning, and our method will not require such access, but we consider such methods for comparison. We also do not assume access to densities or logits from the underlying model. With theses assumptions, our method is applicable on top of API-based models.
We are not aware of prior work on learning algorithms that uses this access model. So, we will compare to a few baselines that either get more access (full finetuning of open source models) or less access (just sampling with different prompts).
@@ 4.2 Theoretical motivation for the Q-probe policy
To motivate the Q-probe policy, it is instructive to consider the limit as we take $k\to\infty$. In particular, we will show that in this limit, the policy converges to the optimal KL constrained policy that maximizes the expected value of the probe $Q_{\theta}$.
@@
[…]
@@
The runtime to get a sample accepted is $M$ iterations in expectation. We can view the softmax version as an approximation of rejection sampling with $k$ in place of $M$. This gives us consistent runtime and parallelization, but does mean that for finite $k$ we are only approximately sampling from the target distribution.
@@ 5.1 Learning from oracle reward feedback
The simplest approach is to simple use mean squared error to learn a $Q$ probe to approximate the oracle reward function directly.
@@
$L_{CE}(\theta)=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{a\sim p_{0}|x}[$ $r(x,a)\log\sigma(Q_{\theta}(x,a)+$ (3)
@@
$\approx\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a\sim p_{0}|x\\ a_{2},\dots,a_{k}\sim p_{0}|x\end{subarray}}\left[-r(x,a)\rho_{\theta}(a,\{a_{i}\}_{i=1}^{k})\right]$ (8)
@@
This PG loss ends up looking much like a contrastive loss, which have traditionally been used for representation learning [Wu et al., 2018, Oord et al., 2018]. Here, the contrastive loss arises naturally since the inference-time procedure of selecting one sample from many requires us to compare and contrast a set of samples. By directly tying the loss to the inference procedure we can force the model to allocate it’s errors in such a way that performs better when selecting a sample by softmax.
@@ 5.2 Learning from preference feedback
The simplest approach to use Q-probes to learn from preferences is use the probe to learn a reward model using a Bradley-Terry model. The per sample loss is:
@@
Finaly, while we did not find it to be useful in practice, it is also possible to parameterize direct policy learning objectives from preference feedback with Q-probes as in DPO [Rafailov et al., 2023]. A full derivation can be found in Appendix B.
@@ 6 Oracle reward experiments
For our first experiments, we evaluate the ability of Q-probes to maximize ground-truth oracle rewards. Specifically, we focus on a program synthesis as a task with oracle rewards given by evaluating test cases. We train probes using the training set from MBPP [Austin et al., 2021] and test on the MBPP test set as well as evaluating generalization to HumanEval [Chen et al., 2021].
@@ 6.1 Setup
We train models on the MBPP train set which consists of 464 programming prompts with test cases. We consider the reward to be 1 if all tests are passed and 0 otherwise. For each training prompt, we can generate as many completions as we want from the base model to automatically label with these rewards. We sample from the base model with temperature 0.8 and top-p 0.95, following [Roziere et al., 2023], unless otherwise noted. For experiments on Code-LLaMA-7B, we take the $26$th hidden layer of the same model for embeddings [^1] Probe performance usually peaks at an intermediate layer [Hewitt and Manning, 2019].. For OpenAI API experiments, we experiment with both embedding API call as well as Code-LLaMA-70B. Unless otherwise stated, the Q-probe is a 1-layer (linear) probe, the optimizer is Adam [Kingma and Ba, 2014], the learning rate is $55e-$ 5, batch size is 1000, and we train for 150 epochs. For the PG loss, we need multiple samples from one prompt to compute the loss. To do this, we group samples by prompt and reshape the batch so it contains 100 problems with 10 samples from each problem.
@@
We consider a variety of baselines. First, we report the average success rate of the base model with default temperature sampling (Baseline (Pass@1)). We also report greedy sampling from the base model (Base-Greedy). We include a few shot baseline where we sample 5 successful completions from the training dataset and put them into context and then sample with temperature 0 (5-shot on successes). We also include a skyline of pass@48 which has oracle access to the ground truth rewards.
For the Code-LLaMA model, we have white-box access to the model so we also add baselines that use LORA finetuning [Hu et al., 2021]. We consider supervised finetuning on the successful completions from the training data followed by greedy decoding (SFT on successes) [Singh et al., 2023, Dong et al., 2023]. We also consider two kinds of rejection sampling alternatives: one using instruction to prompt model to judge its own generation (Prompt RM) and the other using a LORA finetuned reward model instead of a lightweight probe (Finetune RM). At inference time, both rejection sampling baselines adopt hardmax over 48 generations.
@@ 6.2 Code-LLaMA results
We present results for training Q-probes on top of Code-LLaMA-7B in Table 1. The main finding is that Q-probe with the policy gradient loss $L_{PG}$ is the best model. This confirms the idea that finding a loss that is a more direct proxy for the downstream task leads to better outcomes.
At a higher level, it is also important to note the benefits of training such small and lightweight probes. With only a small amount of training we can extract a useful discriminator from the generative model to improve performance.
@@
Finally, Figure 4 attempts to provide some intuition about how $L_{PG}$ differs from $L_{Q}$ and $L_{CE}$ in a way that is beneficial. First and foremost $L_{PG}$ attempts to optimize a proxy of the test metric, expected reward. This experiment tries to look at a lower level to see how this changes the learned models. Intuitively, $L_{Q}$ and $L_{CE}$ treats samples $a$ from the same $x$ independently (since they just sum over all samples), and end up allocating a good amount of capacity to classifying which prompts are hard (causing higher slope in the figure). But the $L_{PG}$ loss forces the model to learn which completions are good when compared to each other for the same prompt. The contrastive nature of this loss helps the model allocate capacity more effectively to the part of the problem that matters: comparing different completions of the same prompt.

### v2

@@ Abstract
We present an approach called Q-probing to adapt a pre-trained language model to maximize a task-specific reward function. At a high level, Q-probing sits between heavier approaches such as finetuning and lighter approaches such as few shot prompting, but can also be combined with either. The idea is to learn a simple linear function on a model’s embedding space that can be used to reweight candidate completions. We theoretically show that this sampling procedure is equivalent to a KL-constrained maximization of the Q-probe as the number of samples increases. To train the Q-probes we consider either reward modeling or a class of novel direct policy learning objectives based on importance-weighted policy gradients. With this technique, we see gains in domains with ground-truth rewards (code generation) as well as implicit rewards defined by preference data, even outperforming finetuning in data-limited regimes. Moreover, a Q-probe can be trained on top of an API since it only assumes access to sampling and embeddings. Code: https://github.com/likenneth/q_probe.
@@ 1 Introduction
On the other hand, there is evidence that the capabilities required for these downstream tasks have already been learned during pre-training, and the task of adaptation is merely to extract them from the wide spectrum of pre-trained capabilities. For example,  Zaken et al. [2021] propose that extremely parameter-efficient finetuning is evidence that the finetuning process is mostly about “exposing knowledge induced by language-modeling training”, while Saunders et al. [2022] find that pre-trained language models are usually better at discriminating than generating answers.
Motivated by this line of thought, we present a lightweight approach to reward maximization. For each downstream task, we keep the whole pre-trained model frozen and only train a small probe which is the same dimension as the residual stream [Alain and Bengio, 2016]. We call our method Q-probe as it “probes” the expected utility of a completion (action) given a certain prompt (state).
To leverage the Q-probe at inference to generate samples, we perform a sort of rejection sampling. Specifically, we first draw $k$ sampled completions from the LLM given the input prompt and also store the embedding of each prompt-completion pair. The Q-probe then predicts a value for each embedding, which determines the logits for the $k$-way softmax distribution that we use to sample the chosen completion. In theory, we show that this procedure maximizes the KL-constrained value of the probe as $k$ tends to infinity.
Figure 1: An illustration of the Q-probe inference procedure. Given a prompt, we use the language model to generate $k=3$ completions (in this case, programs) and the respective embeddings of the $k$ prompt-completion pairs. Then the linear Q-probe maps the features into the logits of a softmax distribution. We obtain our final sample from the Q-probe by sampling from this distribution.
First, we evaluate Q-probes with access to ground truth rewards on coding benchmarks our best Q-probe achieves $17\%$ higher accuracy on MBPP [Austin et al., 2021] compared to the base Code-LLaMA-7B [Roziere et al., 2023] and outperforms finetuning on successes with LORA [Hu et al., 2021] and few shot prompting. Although again, we emphasize that Q-probes are not mutually exclusive with these other techniques and can be combined for even better results. One key component of the results is a novel objective for training the Q-probes via direct policy learning. We find that rather than training the Q-probe to model the rewards, it is more effective to use an importance weighted policy gradient objective. Since we only need access to samples and embeddings we can train Q-probes on API-based models where gains are more modest ($3\%$ improvement over base model) due to a stronger base model and lack of access to internal model embeddings.
@@
Finally, in terms of computational cost, we should note that using a Q-probe requires substantially less training compute, but more inference-time compute when compared to finetuning. In our experiments, we can train a Q-probe in a few seconds (since it is just a 4096-dimensional linear model) whereas even parameter efficient finetuning [Hu et al., 2021] takes several hours. But, at inference, we draw $k$ samples from the base model rather than 1 from a finetuned model, although improvements in parallel and speculative decoding are making batched decoding easier [Fang et al., 2021, Yu et al., 2022, Shen et al., 2024].
@@ 2 Related work
Q-probes leverage the idea of probing to solve reward maximization problems. This idea builds on prior work that uses probes for understanding the internals of neural networks [Alain and Bengio, 2016, Belinkov, 2016, Li et al., 2022]. A probe is a classifier or regressor that takes internal activations of a network as its input and is trained to predict a feature of interest, e.g., part of speech, parse tree depth, or the expected reward in our case.
@@
Rejection sampling for reward maximization is not a new idea. In fact, Gao et al. [2023], Ganguli et al. [2022], Rafailov et al. [2023] also evaluate rejection sampling as one of their baselines. However, their selector model is instantiated by the preference language model trained in a similar way to the first stage of RLHF by Ouyang et al. [2022]. This version of rejection sampling not only involves higher training cost but is also double the inference cost to run the reward model while evaluating the Q-probe is essentially free in comparison to the base model.
@@
An important line of training-free adaptation methods centers around prompting [Salewski et al., 2023] which includes in-context learning (ICL,  Min et al., 2022) and Chain-of-thoughts (CoT,  Wei et al., 2022). Though it enjoys great flexibility, Mosbach et al. [2023] reveal by a closer examination that finetuning still outperforms prompting methods. Prompting could also be sensitive to prompt engineering [Lu et al., 2021] and takes up a valuable context window, limiting the amount of data we can feed into it for a fair comparison with Q-probe and finetuning.
@@
There are also a host of other inference-time techniques designed for coding and reasoning settings [Zhou et al., 2023a, Shinn et al., 2023, Yao et al., 2023]. However, they require access to the feedback from the environment at test time which is different from the one-pass setting considered by us.
@@ 3.1 Feedback: oracle rewards and preferences
Note, there is a large literature of prior work on using reinforcement learning directly to finetune language models when given access to oracle reward functions, e.g., for single turn language tasks [Schulman et al., 2017, Snell et al., 2022, Ramamurthy et al., 2022, Chang et al., 2023] or in multiturn settings [Zhou et al., 2023b, Abdulhai et al., 2023]. In contrast, we focus on a lighter weight approach that only requires training probes, but shows how probe training can approximate traditional RL objectives.
@@
    Online. With online access, we can query the reward or preference of any action $a$ or action pair $(a_{0},a_{1})$ from any context $x_{i}$ in the training set to get $r(x_{i},a)$. This setting is reasonable if we have unit tests for programming or a human in the loop for preference learning.
@@ 3.3 Access to the LLM
    Sampling from the LM distribution $p_{0}$. Given a context $x$ we can sample a completion $a$ from $p_{0}(\cdot|x)$.
@@
We do not in general assume access to the underlying model to allow for finetuning, and our method will not require such access, but we consider such methods for comparison. We also do not assume access to densities or logits from the underlying model. With these assumptions, our method is applicable on top of API-based models.
We are not aware of prior work on learning algorithms that use this access model. So, we will compare to a few baselines that either get more access (full finetuning of open source models) or less access (just sampling with different prompts).
@@ 4.2 Theoretical motivation for the Q-probe policy
To motivate the Q-probe policy, it is instructive to consider the limit as we take $k\to\infty$. In particular, we will show that in this limit, the policy converges to the optimal KL-constrained policy that maximizes the expected value of the probe $Q_{\theta}$.
@@
$\lim_{k\to\infty}\pi_{\theta,k}=\arg\max_{\pi}\operatorname*{\mathbb{E}}_{a\sim\pi|x}[Q_{\theta}(x,a)]-\beta\text{KL}(\pi\|p_{0})$
@@
The runtime to get a sample accepted is $M$ iterations in expectation. We can view the softmax version as an approximation of rejection sampling with $k$ in place of $M$. This gives us consistent runtime and parallelization, but does mean that for finite $k$, we are only approximately sampling from the target distribution.
@@ 5.1 Learning from oracle reward feedback
The simplest approach is to simply use mean squared error to learn a $Q$ probe to approximate the oracle reward function directly.
@@
$L_{CE}(\theta)=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{a\sim p_{0}|x}[$ $r(x,a)\log\sigma(Q_{\theta}(x,a))+$ (3)
@@
$\approx\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a\sim p_{0}|x\\ a_{2},\dots,a_{k}\sim p_{0}|x\end{subarray}}\left[-r(x,a)\rho_{\theta}(a,\{a_{i}\}_{i=1}^{k})\right]$
@@
This PG loss ends up looking much like a contrastive loss, which has traditionally been used for representation learning [Wu et al., 2018, Oord et al., 2018]. Here, the contrastive loss arises naturally since the inference-time procedure of selecting one sample from many requires us to compare and contrast a set of samples. By directly tying the loss to the inference procedure we can force the model to allocate its errors in such a way that performs better when selecting a sample by softmax.
@@ 5.2 Learning from preference feedback
The simplest approach to use Q-probes to learn from preferences is to use the probe to learn a reward model using a Bradley-Terry model. The per sample loss is:
@@
Finally, while we did not find it to be useful in practice, it is also possible to parameterize direct policy learning objectives from preference feedback with Q-probes as in DPO [Rafailov et al., 2023]. A full derivation can be found in Appendix B.
@@ 6 Oracle reward experiments
For our first experiment, we evaluate the ability of Q-probes to maximize ground-truth oracle rewards. Specifically, we focus on a program synthesis as a task with oracle rewards given by evaluating test cases. We train probes using the training set from MBPP [Austin et al., 2021] and test on the MBPP test set as well as evaluating generalization to HumanEval [Chen et al., 2021]. To see if the method generalizes to mathematical capabilities, we carry out experiments on GSM-8K in Subsection 6.4.
@@ 6.1 Setup
We train models on the MBPP train set which consists of 464 programming prompts with test cases. We consider the reward to be 1 if all tests are passed and 0 otherwise. For each training prompt, we can generate as many completions as we want from the base model to automatically label with these rewards. We sample from the base model with temperature 0.8 and top-p 0.95, following [Roziere et al., 2023], unless otherwise noted. For experiments on Code-LLaMA-7B, we take the $26$th hidden layer of the same model for embeddings [^1] Probe performance usually peaks at an intermediate layer [Hewitt and Manning, 2019].. For OpenAI API experiments, we experiment with both embedding API calls as well as Code-LLaMA-70B. Unless otherwise stated, the Q-probe is a 1-layer (linear) probe, the optimizer is Adam [Kingma and Ba, 2014], the learning rate is $55e-$ 5, the batch size is 1000, and we train for 150 epochs. For the PG loss, we need multiple samples from one prompt to compute the loss. To do this, we group samples by prompt and reshape the batch so it contains 100 problems with 10 samples from each problem.
@@
We consider a variety of baselines. First, we report the average success rate of the base model with default temperature sampling (Baseline (Pass@1)). We also report greedy sampling from the base model (Base-Greedy). We include a few-shot baseline where we sample 5 successful completions from the training dataset and put them into context and then sample with temperature 0 (5-shot on successes). We also include a skyline of pass@48 which has oracle access to the ground truth rewards.
For the Code-LLaMA model, we have white-box access to the model so we also add baselines that use LORA finetuning [Hu et al., 2021]. We consider supervised finetuning on the successful completions from the training data followed by greedy decoding (SFT on successes) [Singh et al., 2023, Dong et al., 2023]. We also consider two kinds of rejection sampling alternatives: one using instruction to prompt the model to judge its own generation (Prompt RM) and the other using a LORA finetuned reward model instead of a lightweight probe (Finetune RM). For the latter, we add linear probe to the base policy model at the last residual steam; but different to Q-probe, the whole model is tuned for judging reward with Lora [Hu et al., 2021]. At inference time, both rejection sampling baselines adopt hardmax over 48 generations.
@@ 6.2 Code-LLaMA results
We present results for training Q-probes on top of Code-LLaMA-7B in Table 1. The main finding is that Q-probe with the policy gradient loss $L_{PG}$ is the best model. This confirms the idea that finding a loss that is a more direct proxy for the downstream task leads to better outcomes. The policy gradient loss contrasts many samples for the same prompt, which mirrors the inference procedure and leads to better performance at test time.
Also, recall that Q-probe is easily combined with other methods. To illustrate this, we combine few shot prompting with Q-probe. This leads to even better performance, showing how the different inference procedures are actually leading to complementary improvements in performance that neither approach achieves on its own.
At a higher level, it is also important to note the benefits of training such small and lightweight probes. Because the probe is so small, we can extract a useful discriminator from the generative model with only a small amount of training and use this probe to improve performance.
@@
Finally, Figure 4 attempts to provide some intuition about how $L_{PG}$ differs from $L_{Q}$ and $L_{CE}$ in a way that is beneficial. First and foremost $L_{PG}$ attempts to optimize a proxy of the test metric, expected reward. This experiment tries to look at a lower level to see how this changes the learned models. Intuitively, $L_{Q}$ and $L_{CE}$ treat samples $a$ from the same $x$ independently (since they just sum over all samples) and end up allocating a good amount of capacity to classifying which prompts are hard (causing higher slope in the figure). But the $L_{PG}$ loss forces the model to learn which completions are good when compared to each other for the same prompt. The contrastive nature of this loss helps the model allocate capacity more effectively to the part of the problem that matters: comparing different completions of the same prompt.